### v1

Abstract
Biological intelligence uses a “multiscale competency architecture” (MCA). It exhibits adaptive, goal directed behaviour at all scales, from cells to organs to organisms. In contrast, machine intelligence is only adaptive and goal directed at a high level. Learned policies are passively interpreted using abstractions (e.g. arithmetic) embodied in static interpreters (e.g. x86). Biological intelligence excels at causal learning. Machine intelligence does not. Previous work showed causal learning follows from weak policy optimisation, which is hindered by presupposed abstractions in silico. Here we formalise MCAs as nested “agentic abstraction layers”, to understand how they might learn causes. We show that weak policy optimisation at low levels enables weak policy optimisation at high. This facilitates what we call “multiscale causal learning” and high level goal directed behaviour. We argue that by engineering human abstractions in silico we disconnect high level goal directed behaviour from the low level goal directed behaviour that gave rise to it. This inhibits causal learning, and we speculate this is one reason why human recall would be accompanied by feeling, and in silico recall not.
Keywords:
multiscale competency swarm intelligence consciousness.
1 Introduction
Biological intelligence is sample efficient. It is able to generalise or infer cause after only a few examples. Causal learning facilitates theory of mind, subjective experience and self awareness [bennett2023c, barron2016]. Our purpose here is to better understand how this capability might be engineered. Previous work established that causal learning requires one construct rather than presuppose variables and objects [bennett2023c]. By optimising for policies that express “weak” constraints on behaviour, relations between “contentless” [thompson2016] states are reduced to objects which cause affect [bennett2023b]. These claims were supported by proofs and experiments, but the resulting system is inefficient because it relies upon a high level, static abstraction layer. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. The abstraction layer on which the aforementioned experiments were run conveys an inductive bias. It limits what sort of “causes” can efficiently be learned. How might we achieve efficient causal learning? What sort of abstraction layer would fix this problem?
1.0.1 Biological multiscale competency architecture:
Biology exhibits the capability we desire, so we look to biology for a more suitable abstraction layer. In the 70 years since psychoneural identity theory was proposed [sep-mind-identity], it has become evident that biological intelligence is not neatly confined to a neuronal abstraction layer [ciaunica2023]. Cells are now understood to be an agentic material [ball2023] and biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [morokuma2017], tadpoles with eyes on their backs that see [blackiston2013], and biological “xenobots” that perform simple tasks [levin2020]. Biology uses a “multiscale competency architecture” (MCA), in which lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [fields2020]. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Cancer is an example of what happens when this fails [levin2021], as cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [davies2011]. MCA is relevant to AI because biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [reid2011]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, human intelligence is the high level goal directed behaviour of a collective of cells. If we want to engineer more human-like machine intelligence, then we should mimic an MCA.
1.0.2 Agentic Abstraction Layers:
High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. emphasised the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [ciaunica2023b]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above, like how a CPU based on the x86 instruction set architecture (ISA) would the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts, compelling the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [thompson2007] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate higher level goals.
1.0.3 Summary:
Do MCAs facilitate causal learning? How? Previous AGI research:
  1. 1.
    formalised enactive cognition, to describe goal directed behaviour in terms about which objective claims can be made [bennett2023b, bennett2023a],
  2. 2.
    proved that a weaker policies are optimal in terms of sample efficiency [^1] Also demonstrating a $110-500\%$ improvement in causal learning in experiments. [bennett2023b], causal learning, and
  3. 3.
    argued this facilitates theory of mind, reafference and self awareness [bennett2023c].
Here we:
  1. 1.
    examine the behaviour of MCAs in the context of this formalism, and show weaker policies at higher levels imply weaker policies at lower levels, facilitating what we call multiscale causal learning.
  2. 2.
    We argue a static ISA like x86 hinders causal learning in comparison, and
  3. 3.
    that lower levels “compel” higher. We speculate this may explain why human memory and feeling are intertwined. By disconnecting high level from low in silico, we make it possible to interpret without the impetus of “feelings”.
2 The Formalism
The following definitions are shared with [bennett2024b, bennett2023a]. More extensive explanations of these definitions are given there. We have endeavoured to make key aspects of this paper understandable without the formalism, so readers who wish to should simply skip to the next section. The definitions are here for those who wish to dig into the details.
Definition 1 (environment)
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer)
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless explicitly stated otherwise. If $\mathfrak{v}=P$, then we say that there is no abstraction.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript222e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
Definition 3 ($\mathfrak{v}$-task)
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
2.0.1 Learning and inference definitions:
Inference requires a policy and learning a policy requires a proxy. Strictly speaking, an organism is an embodied policy. However, organisms also adapt to particular tasks, so organisms also have and learn policies.
Definition 4 (inference)
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^3] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
Definition 5 (learning)
  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
By the weakness of a statement, we mean the cardinality of its extension.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$ ($l$ is a cause of $\alpha$). We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child, to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is $<_{w}$.
Definition 6 (uninstantiated-tasks)
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7 (utility of intelligence)
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
(intuitive summary) Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
3 Arguments
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed, and how “the behaviour of subunits percolates up” to higher level goal directed behaviours, which in turn “constrains and facilitates” lower level behaviours [levin2024]. Once that is done we show how an MCA can facilitate more efficient causal learning than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
$cells$$organs$$organism$
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$. The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [bennett2023c] that differentiates observation from intervention in the manner of reafference [barron2016].
$\alpha_{1}$$\alpha_{2}$$\omega$
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2023b, prop. 1,2]. For our cell example, weak policy at the level individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words the construction of a weak policy is not just enabled by an MCA, but is required to form one at all. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; an emergent property we ascribe to the behaviour of the lowest level. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [bennett2023b, proposition 1,2] facilitates causal learning at a single scale [bennett2023c], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 7) for tasks at higher levels, and thus the efficiency of causal learning. That a collective of cells is an “organ” is just an abstract classification we ascribe to the collective’s behaviour. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, an “abstraction layer” is a property we ascribe to the emergent behaviour of many parts interacting. It appears to us as a second order system necessarily implied by the outputs $\mathfrak{v}$-tasks at different scales, because we see the whole rather than the parts. In the above example, $\omega$ is just $\alpha_{1}$ and $\alpha_{2}$ together, but to understand how we might “intervene” in the causal chain and disconnect emergent properties from the system which gave rise to them, we need to abstract away the individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$. We use this fact to justify representing policies abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour which has emergent properties. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Proposition 1 (propagation)
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
Proof
If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies). We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$. If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$. ∎
When we say we want efficient causal learning in computers, what we’re saying is that we want to mimic the adaptability of high level goal directed behaviour in humans. Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
$cells$$organs$$human$$abstractions$
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.
$human$$ISA$$abstractions$
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
Dropping Hume’s Guillotine: Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just abstractions we ascribe to the behaviour of the lowest level collective). Previous work argues learned behaviour must be motivated by attraction and repulsion from aspects of the environment [bennett2023c]. This eventually culminates in “feelings” at higher levels [bennett2023d]. Assuming that is true, then by intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. We suggest that a computer is not a “syntactic engine”, as it is sometimes portrayed [kim2011]. Syntax has meaning to the computer, otherwise it would do nothing. However the “meaning” of an instruction is what it does, just as the “meaning” of a declarative program is its truth conditions. What a computer is, is meanings realised in circuitry by human hands. That is why it “interprets” syntax as it does. We speculate one reason computers may not “feel” and “experience” meaning as a human would is because there is a static, hard coded abstraction layer that allows high level goal directed behaviours to exist in the absence of the low level goal directed behaviours of which they are an emergent property. As a result, interpretation is fixed rather than fluid and motivated (e.g. by affect).
[…]