### v2

# Multiscale Causal Learning

Abstract
Biological intelligence is more sample-efficient than artificial intelligence (AI), learning from fewer examples. Here we answer why. Given data, there can be many policies which seem “correct” because they perfectly fit the data. However, only one correct policy could have actually caused the data. Sample-efficiency requires a means of discerning which. Previous work showed sample efficiency is maximised by weak-policy-optimisation (WPO); preferring policies that more weakly constrain what is considered to be correct, given finite resources. Biology’s sample-efficiency demonstrates it is better at WPO. To understand how, we formalise the “multiscale-competency-architecture” (MCA) observed in biological systems, as a sequence of nested “agentic-abstraction-layers”. We show that WPO at low levels enables synthesis of weaker policies at high. We call this “multiscale-causal-learning”, and argue this is how we might construct more scale-able, sample-efficient and reliable AI. Furthermore, a sufficiently weak policy at low levels is a precondition of collective policy at higher levels. The higher level “identity” of the collective is lost if lower levels use an insufficiently weak policy (e.g. cells may become isolated from the collective informational structure and revert to primitive behaviour). This has implications for biology, machine learning, AI-safety, and philosophy.
keywords:
multiscale competency, AI, causal learning, sample efficiency
1 Introduction
Intelligence is often understood as predicting the outputs of a system, given inputs [1, 2, 3]. In that frame, learning is the act of leveraging data [^1] A set of inputs and outputs known to correspond to a particular system. We also refer to this as training or historical data. to improve predictive accuracy. It is an optimisation problem. The cause of data is the policy [^2] By policy, we just mean a constraint on outputs and inputs. In other words, a policy is a declarative definition of behaviour. There is an equivalence between imperative and declarative programs [4] so intuitively, consider a policy to be a program that dictates behaviour. We use the word policy for the sake of consistency with related work and to avoid having to use two different terms for the same thing. We do not mean only the more narrow definition of policy used in reinforcement learning. which originally generated it. To borrow a philosophical term [5], it is the policy which was the original “causal intermediary” between outputs and inputs. In machine learning, training data are sets of examples of goal directed behaviour; an ostensive definition of how to complete a task. A learning algorithm synthesises a policy which “fits” training data. Intuitively, any policy that fits is “correct” for that training data. However, the fact that a policy is correct with respect to data is not enough information to determine whether that policy was the cause of that data. Data could be the result of any number of policies, and if one wishes to accurately predict outputs given inputs that aren’t in the training data, then one must identify that policy which originally generated the data. The true cause. The closer one’s policy is to the cause, the more accurately one will be able to predict outputs given new inputs one hasn’t seen before. Conversely, the better one is at identifying the likely causes of data, the less data one will require to learn how to complete the task. The less data one requires, the more “sample efficiently” one learns.
1.1 Active Inference is Not Enough
Active inference [6, 7] formalises how organisms learn and adapt. It describes living systems in terms of a machine learning approach called variational Bayes. Like other machine learning approaches, this treats learning as a process which synthesises a policy which happens fit data. A policy that is correct relative to historical data, but may not be the true cause of that data. Yet biological intelligence (BI) is far more sample efficient than artificial intelligence (AI). Organoids learn far quicker than deep reinforcement learning agents in silico [8]. Yes, a robust AI will eventually learn cause and effect [9], but for now biology remains for more efficient. How is it that BI is so much more sample efficient than AI? This suggests there is more to the picture than active inference. BI is not just synthesising a policy which “fits” historical data; it is going a step further to figure out which among correct policies are likely to have caused historical data. How?
1.2 Weak-Policy-Optimisation is Causal Learning
Previous work already demonstrates how optimal sample efficiency [10] and causal learning [11] are achievable. Every policy implies a constraint on possible outputs and inputs; a constraint on how many possible tasks [^3] A task means a set of inputs and a corresponding set of outputs. A task is an ostensive definition of a goal or purpose. This is formally defined later in the paper. will be completed by this policy. A more specific policy completes fewer possible tasks, while a policy that implies a weaker constraint completes more possible tasks [12]. Past research into causal learning and sample efficiency found that the correct policies most likely to have caused historical data imply the weakest [^4] Formal definitions for policy weakness are given in the next section. constraint [10, 11] [^5] See Bennett’s Razor and related proofs in [10] for sample efficiency, and [11] for causality arguments.. Experiments demonstrated that a preference for weak policies correctly identifies the cause of historical data at $110-500\%$ the rate of a complexity based induction system [10]. Gradient descent does not optimise for weak policies, which is one reason why deep learning is so sample inefficient [13]. These results are further supported by the contravariance principle proposed by Cao and Yamins [14]. They argue tasks that more tightly constrain the space of correct policies [^6] A larger, more varied set of inputs and outputs implies a smaller set of correct policies. make it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. Given the sample efficiency exhibited by organoids [8], it would appear biology is far better at optimising for weak policies than current approaches to machine learning. Causal learning via weak-policy-optimisation [^7] Meaning learning using the weakness proxy as formally defined in the next section and in [3] (WPO) is clearly possible in silico, but existing implementations [10] are computationally expensive. One goal of our research is to better understand how the efficiency we observe in biology might be engineered.
1.2.1 How Weak Policy Optimisation Constructs Causal Identities
Previous work also established that causal learning requires one construct rather than presuppose variables and objects [11]. Causal identities are learned by constructing $X$ such that $X$ causes $Y$, not by presupposing the environment is divided into objects including $X$ and then intervening to test if $X$ causes $Y$ [15]. By optimising for policies that express “weak” constraints on behaviour, otherwise meaningless primitives888e.g. Declarative programs the truth value of which can be known, but for the truth conditions can never be known because environmental states are observable only partially and indirectly [3, 16, 15]. or very low level abstractions may be clustered together to form “causal identities” for higher level abstractions, like variables [11]. The “do” operator which is used to distinguish intervention from observation in causal inference problems can be substituted by introducing variables for interventions [17]. That is why one can learn cause and effect via WPO [11], by constructing causal identities for interventions.
1.2.2 How Causal Identities Relate to Abstraction Layers
The construction of causal identities is a process by which an environment can be broken up into abstract classifications that are inherently meaningful (motivated and relevant to a goal). In philosophical terms, if one is motivated by attraction and repulsion from “contentless” [18] states [^9] Contentless meaning categorised, value-less, monolithic states with no discernible internal structure., then WPO is a means by which those states can be reduced to objects or events which cause the affect which motivates an organism [15]. In computational terms, this amounts to constructing an abstraction layer that is most fit for a purpose [3, 16]. For example, the x86 instruction set architecture (ISA) has hard coded interpretations of certain sorts, such as numbers and arithmetic operations [^10] These amount to policies by the formal definition we use [3, 16]. As a result, the ISA can more efficiently address some problems than others. It is for this reason field programmable gate arrays exist as an alternative to static ISAs like x86; so that specific architectures can be constructed to more efficiently address specific problems.
1.2.3 Static Abstraction Layers and Computational Dualism
The ability to construct causal identities is easily hindered by the presupposition of static abstractions [11, 3]. This is because one cannot construct weaker abstractions if one is forced to use a static set of high level abstractions, like numbers and arithmetic operations. This is symptomatic of problem called “computational dualism[3]. Where 16th century Cartesian interactionist dualism proposed separate mental and physical substances interacting through the pineal gland, computational dualism is the concept of software distinct from hardware. For example, Hinton described software as “immortal”, existing infinitely many times across devices [19]. According to the argument against computational dualism [3], immortal computations do not exist. There are only finitely many instances of mortal computations. Intuitively, when one loads software onto a computer, no object moves into the computer. All that happens is that the state of hardware changes. Addressing computational dualism revealed weakness to be the confounding variable that correlates simplicity with generalisation [16, Prop. 2], allowing objective claims to be made regarding theoretical intelligent agents [3], addressing the subjectivity problems of complexity based induction systems in interactive settings [20]. This is all relevant to us because static abstraction layers are based on computational dualism. They seem to suggest software is something other than the state of hardware. Something interchangeable. Move-able. This means static abstraction layers are necessarily one size fits all, rather than optimised for particular purpose. The inefficiency introduced at the abstraction layer affects everything built within it, including any software based machine learning systems. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias. Presupposing the world is formed of numbers, objects and operations of a specific sort limits which cause and effect relations can efficiently be learned. The sample efficiency exhibited by biology suggests it uses a more versatile alternative to static abstraction layers based on computation dualism; one more suitable for WPO.
1.3 Multiscale Competency Architecture
Biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [21], tadpoles with eyes on their backs that see [22], and biological “xenobots” that perform simple tasks [23]. As biology becomes more tractable, the similarities between biology and human artifice (e.g. computers) have become more obvious. Cells are now understood to be an agentic material [24] for the purpose of engineering. They are but one scale in the “multiscale competency architecture” (MCA) of biology. In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [25]. Adaptation takes place at every scale. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [26]. Cancer happens when this fails [27]. Cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [28]. Biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [29]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, there is a great deal of evidence [26] to suggest that human intelligence is nothing more than the high level goal directed behaviour of a collective of cells. If we want sample-efficient AI, then we should mimic an MCA.
2 Materials and Methods
To summarise what has been covered so far:
  1. 1.
    Our goal is to understand why BI is so much more sample efficient (able not just to synthesise a policy that “fits” data, but to infer which among such correct seeming policies originally generated or caused the data).
  2. 2.
    As sample efficiency is maximised by WPO, our purpose is to understand how biology is able to synthesise weak policies so effectively.
  3. 3.
    A promising avenue appears to be the MCA [25] used by biological systems, instead of the static abstraction layers used in conventional computers (which exhibit the shortcomings of computational dualism [3]).
Hypothesis: The agentic, nested structure of MCAs allows more effective and efficient WPO than would be possible within a static layer like x86.
2.1 Agentic Abstraction Layers
To begin to understand how this might be possible, we must formalise MCAs in rigorous, mathematical terms. High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [30]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above; like how a CPU based on the x86 instruction set architecture (ISA) interprets the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts. It compels the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [31] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
2.2 Definitions
We begin with a formalism [3] of enactive cognition [31]. We will use it to formalise MCAs in the next section. More extensive explanations of these definitions are given elsewhere [16, 3], and we have endeavoured to make key aspects of this paper understandable without the formalism; readers who wish to should simply skip to the next section.
Definition 1. [environment].
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
Definition 2. [abstraction layer] .
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless stated otherwise [^11] This is because a bounded system can contain only a finite amount of information [32] in a spatially extended environment like our own [16].. If $\mathfrak{v}=P$, then we say that there is no abstraction.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript121212e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
Definition 3. [$\mathfrak{v}$-task] For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2.2].
  • A correct output is a correct completion of an input.
2.2.1 Learning and inference definitions:
Inference requires a policy and learning a policy requires a proxy. Strictly speaking, an organism is an embodied policy. However, organisms also adapt to particular tasks, so organisms also have and learn policies.
Definition 4. [inference].
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^13] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
Definition 5. [learning] .
  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
By the weakness of a statement, we mean the cardinality of its extension.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$ ($l$ is a cause of $\alpha$). We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [10]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child, to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is $<_{w}$.
Definition 6. [uninstantiated-tasks] The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7. [utility of intelligence] Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
(intuitive summary) Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
3 Arguments and Results
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed; how “the behaviour of subunits percolates up” to determine higher level behaviour, which in turn “constrains and facilitates” lower level behaviours [26]. Once that is done we show how an MCA can facilitate more efficient WPO than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
$cells$$organs$$organism$
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$ (meaning there the number of possible configurations is finite). The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [11] that differentiates observation from intervention in the manner of reafference [33].
$\alpha_{1}$$\alpha_{2}$$\omega$
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy (what we have been referring to as WPO). This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [10, prop. 1,2]. For our cell example, a sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words WPO is not just facilitated by an MCA, but required for high level behaviour. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; something we ascribe to the behaviour of the lowest level. For the sake of intuition; different levels are patterns within one larger pattern, and we simplify the larger pattern by dividing it up. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [10, proposition 1,2] facilitates causal learning at a single scale [11], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 2.2.1) for tasks at higher levels, meaning reduced capacity for WPO. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, x86 is a classification we ascribe to the product of human behaviour. It is a higher level behaviour of an MCA that includes humans. To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels. To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions). We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Theorem 1.
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
Proof of Theorem 1.
If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies). We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$. If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$. ∎
Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
$cells$$organs$$human$$abstractions$
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. No graceful degradation. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.
$human$$ISA$$abstractions$
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [16], because bounded systems can contain only finite information [32]. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
3.1 Precondition for Collective Identity
According to Cao and Yamins [14] tasks are more “difficult” when they have fewer correct policies, and “easier” when they have many correct policies. In our formalism, Cao and Yamins’ position implies $\mathfrak{v}$-task $\alpha$ is “easier” than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. Collectives that exhibit high level goal directed behaviour that is “easier” have more correct high level policies to choose from. If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder”141414i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$. If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, and if the system is robust and exhibits graceful degradation, then part of a collective may “break off” (loosening the constraint on correct policies, for both the breakaway part and the part left behind). In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to “selective forgetting” [11] to cope with false negatives, but instead of forgetting data inconsistent with an otherwise good policy, we would “discard” a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
This is analogous to cancer as cells isolated from a collective informational structure, that lose their high level identity, and pursue lower level goals [27, 28]. If $\alpha$ is a cell, it can now revert to primitive transcriptional behaviours and act like cancer.
4 Discussion
We have formalised MCAs and shown the capacity for WPO at high levels implies WPO at low levels. Causal identities are constructed at different scales. Hence we call it multiscale-causal-learning (MCL). This seems to indicate MCA-like designs could result in more efficient and reliable AI in future.
4.1 Dropping Hume’s Guillotine
Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole). Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [11]. By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. This relates to the hard problem of consciousness [34]. According to our formalism, all abstractions constructed by living things are causal identities “goal directed” in the sense of being constructed in response to motivating affect. In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them. It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements) [15]. Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
4.2 Future Research
4.2.1 AI-Safety, Distributed and Cyberphysical-Systems
The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [35]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [36]. This involves constructing advanced general AI of level ASL-4 [37] [^15] Meaning an artificial general intelligence [38] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI. in a cyberphysical system. That system seems intended to safeguard against threats from both within and without. The idea of a “rogue AI” from within this system is analogues to cancer, adopting an “unsafe” policy outside the scope of what is permitted by collective. This suggests we could formalise such a system as an MCA, in order to prove properties relevant to AI safety. Control of cyberphysical and distributed systems in general might be framed in terms of an MCA, in order to better understand how different levels in a hierarchy or different parts of a horizontal system constrain each-other.
4.2.2 Active Inference
The sample efficiency of organoids [8] reveals what active inference misses; WPO. Future research might integrate WPO with active inference. Surprise minimisation [7] could be considered an extremely weak policy in and of itself. However, what really matters is fitness [15], not surprise. In adopting machine learning as its foundation, active inference may have inadvertently dropped Hume’s Guillotine, emphasising the wrong goal.