### v3

# Are Biological Systems More Intelligent Than Artificial Intelligence?

Abstract
Is a biological self-organising system more ‘intelligent’ than an artificial intelligence? If so, why? We frame intelligence as adaptability, and explore this question using a mathematical formalism of enactive causal learning. We extend it to formalise the multilayer, multiscale, bottom-up distributed computational architecture of biological self-organisation. We then show that this architecture allows for more efficient adaptation than the static top-down interpreters typically used in computers. To put it provocatively, biology is more intelligent because cells adapt to provide a helpful inductive bias, and static interpreters do not. We call this multilayer-causal-learning. However it inherits a flaw of biological self-organisation. Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour. We show that, in the context of our formalism, failure states like cancer occur when systems are too tightly constrained by the abstraction layer in which they exist. This suggests control should be distributed (bottom-up rather than top-down) to ensure graceful degradation. We speculate about what this implies for systems in general, from machine learning hardware to human organisational and economic systems. Our result shows how we can design more robust systems and, though theoretical in nature, it lays a foundation for future empirical research.
1 Introduction
Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7]. They interact with their environment to achieve homeostatic and reproductive goals. Some proactively learn [8], storing information about the past in order to anticipate future needs. To make accurate predictions in environments of sufficient complexity, a system must be able to learn cause and effect [9, 10, 11]. For our purposes we consider the ‘intelligence’ of a system to be a measure of the resources it needs to correctly identify cause and effect [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper. Ours is primarily informed by [12], which is adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. The more efficient the learning system, the wider the range of environments in which it can satisfy goals. It is not possession of skills, but an ability to acquire skills with limited resources. Being an embodied formalism, this also accounts of energy and other resources consumed.. It is a measure of how efficiently one adapts [^2] Primarily sample and energy efficiency. A ’sample’ is information used to learn. When a teacher provides an example of a math problem being solved, they are providing a sample. If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B..
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) of those that learn [17, 18, 19, 20]. Contemporary approaches to machine learning are notoriously inefficient [21]. They need vast quantities of data and energy learn tasks most humans can pick up after one or two examples. Even organoids, small collectives of cells grown in a lab, appear to be more ’intelligent’ in this sense than contemporary machine learning [22]. Why is this the case? Are biological systems more intelligent, or is something else going on?
Some suggest that the relative inefficiency of ML may be down to the fact that biological systems are active inference systems, and machine learning systems are typically not [23, 6, 24]. Active inference is much more efficient [25, 26, 27], but it may only be part of the picture. As of 2024 active inference systems [^3] In particular the models submitted by Jack Cole and his employer, Minds AI. applied in the context of the ARC benchmark [14] [^4] For additional information, see https://arcprize.org for sample efficient adaptation still under-perform humans by a substantial margin [^5] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and which is disqualified from ARC because it does not meet the computational efficiency requirements.. Others hold that the reason biological systems appear to be more intelligent is because of an ‘inductive bias’ [14]. This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to efficient adaptation. If an inductive bias were the only reason for the discrepancy between biological self-organisation and machine learning, that would suggest biological systems are not more intelligent than artificial intelligence, but more knowledgeable.
To reiterate, we seek to explain why biological self organisation seems to be more sample (and energy) efficient than contemporary machine learning. Is there something biological systems do that allows them to be more ’intelligent’ than artificial intelligence in its contemporary form? We suggest the answer lies in how biological self organisation distributes control at a different scales, and levels of abstraction. To do so we must first give some background on causal learning which informs our hypothesis.
1.1 Causality
Cause and effect are typically represented as an acyclic graph.
causeeffect
Complex environments often complicate the ability to identify causes. For example, confounded events may seem to cause one another when in fact they are merely correlated. This can lead one to draw erroneous conclusions about cause and effect.
not the causeeffectconfounder
By actively intervening in the environment a system can experiment to isolate causes. To do so, a system must be able to differentiate between events it passively observes and events it has caused [10, 9, 28]. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relation between them. This works well in theory for artificial intelligence [29], but requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts). This needs only attraction and repulsion from physical states [30], otherwise known as valence. We take this latter approach.
We can frame causal learning in computational terms. Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence. Assume we have fed a set of inputs into a computer, and for each input the computer responded with a particular output. The computer is running a ’policy’ program which maps inputs to outputs [^6] We’ll call all such programs policies for the sake of this paper. This should not be confused with the more specific notion of a policy from reinforcement learning [19].. That policy is the cause of the input-output history we’ve observed so far. We want to know the policy that caused historical outputs so that we can use it to predict future outputs. The problem is that more than one policy exists which could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause. If we wish to know which policy actually caused the outputs, we need way to narrow down the possibilities even further.
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [31]. One way is to can compare their form, and choose programs which are simpler. However, in an interactive setting simplicity is an unreliable indicator of plausibility [32, 31]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data. Adaptive systems frequently satisfy the same functional constraints by different means [33], so we can compare different sorts by the constraints they satisfy. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the greater the variety of future behaviours it permits. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16]. In theoretical terms the policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [34]. This has been used to establish local upper bounds on embodied ‘intelligence’ [12].
ML systems do not typically optimise for either weaker constraints on function, or simplicity of form. This at least partially explains their relative sample inefficiency [35, 34]. However, even when ML systems are modified to optimise for weakness, their capacity to do so limited by the “abstraction layer”. An abstraction layer is like a causal isolator [36] between a software “mind” and its environment. For example, the Python programming language is interpreted by the Python interpreter. In the end, Python gets translated into machine code, which hardware like an Apple M2 system on a chip then interprets to physical effect based on the ARM instruction set architecture. An abstraction layer is just physical embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ’intelligence’ [12]. In other words, physical embodiment imposes severe limitations, and the behaviour of software is determined entirely by the hardware and software abstraction layers in which it runs. Intelligence is primarily a matter hardware, not software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [37, 38, 39].
Put another way an embodied policy is a physical structure, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. A more specific tool completes fewer possible tasks, while a tool that implies a weaker constraint completes more possible tasks [21]. By preferring weaker policies, one avoids unnecessary assumptions and so adapts as efficiently as possible, meaning the weakest policies complete a ‘wider variety of tasks’ than all others [^7] As an informal epistemological razor, this means “explanations should be no more specific than necessary”. Intuitively, this is a similar idea to the principle of maximum entropy [40, 41], but it accounts for the subjectivity of complexity and the teleological nature of language [31]. For further intuition, a corollary to the idea that a weaker policy is more likely to be the cause of data is the inventors paradox, or equivalently the contravariance principle proposed by Cao and Yamins [42]. They argue tasks that more tightly constrain the space of correct policies make better training sets. A larger, more varied set of inputs and outputs implies a smaller set of correct policies. That makes it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. See proofs 1 and 2 for more details.. To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
That biological self-organising systems adapt efficiently suggests they are able to embody weaker policies than today’s ML systems. They attain higher local upper bounds on embodied intelligence [12]. This suggests they have a means of optimising not just the choice of policy, but the embodied abstraction layer in which that policy exists.
In a typical computer, the abstraction layers are static. For example, a Python interpreter does not learn and adapt to suit the program it is running. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias. When we say that biological systems can optimise the abstraction layer, this is like saying they can “move” this window to learn a different inductive bias. The question is, what is the abstraction layer in the context of a biological self-organising system?
We suggest the biological equivalent of abstraction layers are the layers of a “multiscale competency architecture” (MCA). In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [43]. Adaptation takes place at every scale, and cells are like an agentic material [44]. Cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [45]. Each cell pursues goals, and then ever larger groups of cells pursue collective goals at ever larger scales. Biological collectives exhibit high level goal directed behaviours beyond the capability of the individual parts [46]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In computational terms this is a distributed system, like a von Neumann architecture [33].
Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than artificial intelligence because their distributed, multilayer, multiscale architecture allows for higher local upper bounds on ’intelligence’. This is because it can shift abstraction layers to convey helpful inductive biases. Our results, though purely theoretical, lay a foundation for future empirical research.
In section 2 introduce the formalism. In section 3 then argue this formalism represents an MCA and explain how. In the context of that formalism, we will formally show that an MCA allows for higher local upper bounds on embodied ‘intelligence’. We call this multilayer-causal-learning (MCL). In section 4 we explore how MCL may inherit the flaws of biological self-organising systems as well as their increased adaptability. We formalise a failure state analogous to cancer, and show that this state is more likely when a system is more tightly constrained. Finally, we relate these results to systems in general. We speculate about the possible implications, from more efficient hardware for machine learning to improved human organisational systems.
2 Formalism and Proofs
As part of our argument we attempt to describe MCAs in mathematically precise terms. Recent research has stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [47]. Taking our cue from that approach, we will formalise MCAs as a series of nested, adaptive, agentic abstraction layers. Enactivism frames cognition as an interaction between a self-organising system and its environment environment [48, 49, 50]. We merely extend this process inwards, to formalise the interactions within the system itself. A non-agentic, static abstraction layer passively interprets the abstraction layers above. By an agentic abstraction layer, we mean one that has its own agenda. It adapts and compels the layers above to serve its own ends, while they each serve their own increasingly specific goals within the bounds of what lower levels permit. Higher levels are compelled by lower to learn a policy “aligned” with lower levels. Lower levels change their interpretation to enact [48] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
Except for the last definition, all the of the definitions that follow are from preceding work [12, 31]. They formalise self-organisation in terms of constraints, like boundary conditions, as this lends itself to enactive depictions of intelligence [51]. We have endeavoured to make key aspects of this paper understandable without the formalism. Readers who wish to should simply skip to the next section.
{adjustwidth}
11.6pt11.6pt
Definition 1 (environment).
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised [^8] Realised meaning it is made real, or brought into existence. by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer).
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript999e.g. $E_{l}$ is the extension of $l$..
$L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table. Intuitively a sensorimotor system is an abstraction layer. Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
{adjustwidth}
11.6pt11.6pt
Definition 3 ($\mathfrak{v}$-task).
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
2.1 Learning and inference definitions
Inference requires a policy and learning a policy requires a proxy. Strictly speaking, an organism is an embodied policy. However, organisms also adapt to particular tasks, so organisms also have and learn policies.
{adjustwidth}
11.6pt11.6pt
Definition 4 (inference).
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^10] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
{adjustwidth}
11.6pt11.6pt
Definition 5 (learning).
A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$. We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [34]).
Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns (it chooses policies that complete a wider variety of tasks, and thus are more sample and energy efficient choices), the more efficient the proxy. The most efficient proxy is weakness (see proofs 1 and 2, or [34]), which is why we’re using it here.
{adjustwidth}
11.6pt11.6pt
Definition 6 (uninstantiated-tasks).
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7 (utility of intelligence).
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
Uninstantiated tasks let us represent the same task in different abstraction layers. We can then use utility to measure how “good” an abstraction layer is for a particular task [12]. This recognises that abstraction layers are not value-neutral. Each is better or worse for some tasks. Only the absence of abstraction is entirely neutral.
2.2 Formalising Multilayer, Multiscale Architectures
The following definition is explained at length over the remainder of this paper. Refer back here as needed.
{adjustwidth}
11.6pt11.6pt
Definition 8 (multiscale competency architecture).
  • The causal chain of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$.
  • $\mathfrak{f}:2^{L_{\mathfrak{v}}}\rightarrow 2^{P}$ is a function that takes the extension $E_{\pi}$ of a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
  • The state of the MCA is a sequence of policies $\langle\pi^{0},\pi^{1}...\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1}...\mathfrak{v}^{n}\rangle$ such that $\mathfrak{v}^{i+1}=\mathfrak{f}(E_{\pi^{i}})$ and $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v^{i}})}$.
In the absence of abstraction where the system is seen as nothing more than the sum of its parts, the MCA is just a task $\lambda^{0}(\mathfrak{v}^{0})$, allowing us to look at the system across scales of distribution. We say the MCA is over-constrained when there exists $i<n$ s.t. and $\Pi_{\lambda^{i}(\mathfrak{v^{i}})}=\emptyset$, and multilayer-causal-learning (MCL) occurs when the MCA is under-constrained and the proxy for learning is weakness. Note that the vocabulary is different at each level in the MCA, which means each has its own generational hierarchy of tasks. By a higher level of abstraction in the MCA, we mean a later task in the causal chain.
2.3 Proofs
Here we present three proofs. The first two demonstrate the optimality of the weakness proxy, and the last show that adaptability later in the causal chain necessitates adaptability earlier in the chain. Preliminary proof and experimental results showing weakness to be the optimal proxy were published in conference proceedings [34]. However, the proofs contained minor errors. Most notably the earlier proof of sufficiency failed to account for task equivalency and so miscounted the total number of tasks, rather than counting the total number of equivalence classes [^11] Credit goes to Nora Belrose for spotting this error.. This does not change the result, but for the sake of rigour these errors must be addressed. To that end, here we present more detailed alternatives to those proofs, which once again show the optimality of weakness as a proxy.
Proposition 1 (sufficiency).
Assume $\alpha\sqsubset\omega$. The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^12] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
Proof.
You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\pi\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\pi\in\Pi_{\omega}$:
  1. 1.
    For every $\pi\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\pi}\in\Gamma_{\mathfrak{v}}$ s.t. $O_{\gamma_{\pi}}=E_{\pi}$, meaning $\pi$ permits only correct outputs for that task regardless of input. We’ll call the highest level task $\gamma_{\pi}$ s.t. $O_{\gamma_{\pi}}=E_{\pi}$ the policy task of $\pi$.
  2. 2.
    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
  3. 3.
    If a policy $\pi$ is correct for a parent of $\omega$, then it is also correct for $\omega$. Hence we should choose $\pi$ that has a policy task with the largest number of children. As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\pi}$ or a child thereof.
  4. 4.
    For the purpose of this proof, we say one task is equivalent [^13] This is because switching from $\beta$ to $\zeta$ s.t. $I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances. This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs. $O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
  5. 5.
    No two policies in $\Pi_{\alpha}$ have the same policy task [^14] Every policy task for policies of $\alpha$ is non-equivalent from the others.. This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
  6. 6.
    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^15] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
  7. 7.
    For any given $\pi\in{\Pi}_{\alpha}$, the extension $E_{\pi}$ of $\pi$ is the set of outputs $\pi$ implies. The subset of $E_{\pi}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\pi}$ [^16] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
  8. 8.
    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|$ increases monotonically with $\left|E_{\pi}\right|$, because for all $e\in E_{\pi}$, $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$. This is because $L_{\mathfrak{v}}=\overline{E_{{I}_{\alpha}}}\cup E_{{I}_{\alpha}}$ and for all $\pi\in\Pi_{\alpha}$, $E_{\pi}\subseteq L_{\mathfrak{v}}$.
  9. 9.
    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\pi$ generalises. It increases monotonically with the weakness of $\pi$.
  10. 10.
    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\pi\in{\Pi}_{\alpha}$ generalises to $\omega$ is
    $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\pi}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
$p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\pi}\right|$ is maximised. Recall from definition 4 that $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$. $\pi$ that maximises $<_{w}$ will also maximise $p(\pi\in{\Pi}_{\omega}\mid\pi\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$. Hence the weakness proxy maximises the probability that [^17] Subsequently it also maximises the efficiency with which a parent $\omega$ is learned from a child $\alpha$. a parent $\omega$ is learned from a child $\alpha$. ∎
Proposition 2 (necessity).
To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
Proof.
Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
  1. 1.
    If $\pi\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\pi}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\pi}$.
  2. 2.
    If $\left|E_{\pi}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\pi}$.
  3. 3.
    Therefore generalisation is only possible if $\left|E_{\pi}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
  4. 4.
    For any two hypotheses $\pi_{1}$ and $\pi_{2}$, if $\mid E_{\pi_{1}}\mid<\mid E_{\pi_{2}}\mid$ then the probability $p(\left|E_{\pi_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\pi_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
  5. 5.
    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised. To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
To select the weakest hypothesis, it is necessary to use the weakness proxy. ∎
Proposition 3.
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be.
Proof.
If $\mathfrak{a}\subset\mathfrak{b}$ then $\epsilon(\lambda^{i+1}(\mathfrak{a}))<\epsilon(\lambda^{i+1}(\mathfrak{b}))$, meaning if $\mathfrak{b}$ is the vocabulary at $i+1$, then it will be possible to construct weaker policies than if $\mathfrak{a}$ is the vocabulary (intuitively, a larger vocabulary enables a wider range of policies). We consider two policies $\pi^{i}_{\mathfrak{a}}$ and $\pi^{i}_{\mathfrak{b}}$ which could be the policy $\pi^{i}$ at $i$. If $\mathfrak{a}=\mathfrak{f}(E_{\pi^{i}_{\mathfrak{a}}})\subset\mathfrak{f}(E_{\pi^{i}_{\mathfrak{b}}})=\mathfrak{b}$, then $\pi^{i}_{\mathfrak{a}}<_{w}\pi^{i}_{\mathfrak{b}}$, meaning an enlarged vocabulary at $i+1$ implies a weaker policy at $i$. ∎
3 Multiscale, Multilayer Causal Learning
To begin, we must formally describe the architecture before we can consider its strengths and weaknesses in comparison to a typical computer. Hence we describe how $\mathfrak{v}$-tasks can describe the multiple scales and layers of abstraction in a biological system. We frame this first by looking at distribution of control within the system, and second by formalising the levels of abstraction at which we can view the system as a whole. Intuitively, this means we formalise different cross sections of the system.
Looking at how control is distributed across different scales within the system we can see how the individual parts that make up a collective each possess their own policy or “identity”, and how that will (or will not) align with a shared collective policy or identity. This treats the system as nothing more than the sum of its parts. It is about self organisation, and how control is distributed within the resulting system. For example, individual neurons in the brain engage in predictive processing of eachother, and emulating this approach has yielded promising results in artificial neural networks [25, 26, 27]. They form a network with a single coherent policy, and yet at the same time are individual neurons engaging in active inference.
In contrast abstraction is about how the behaviour of cells enact an organ, and the behaviour of organs enacts an organism which enacts even more abstract behaviours like, for example, language. Each of these is a layer of abstraction. Layers exist in a causal ‘chain’ (a directed acyclic graph) where each layer (node) is the goal directed behaviour of the layer that preceded it. For example, cells form organs which form organisms [45].
$cells$$organs$$organism$
3.1 Distributing Self Organisation Across Scales
The formalism we extend is enactive [12, 30], describing a self-organising system by describing what it does, rather than what it is. An embodied, embedded system can only do so many things, and so its behaviour is expressed using a finite vocabulary $\mathfrak{v}$ of declarative programs. These programs describe everything the system can do. For example if a system can be in a state of pain, then there are programs in $\mathfrak{v}$ which are “true” when the system is in such a state. Since states of the environment are mutually exclusive, $\mathfrak{v}$ implies an embodied formal language. By language here, we do not mean like the words humans use to communicate. We mean a language of concrete computations [52], where every statement in that language is rather than describes part of the state of a system (put another way, the interpreter is not assumed to be a UTM, but physical laws). For example, consider the language implied by a human being. If the human raises their arm, then they are expressing a statement in their embodied language. That statement then constrains what the body can do next, because everything that happens next begins from the point where the person had their arm raised [30].
We frame self organisation in the same terms as the computer analogy we employed in the introduction. The behaviour of a self-organising system is goal directed, meaning there exists a constraint on its behaviour that defines that goal. The behaviour that satisfies the goal is expressed in the embodied formal language; it is a set of inputs and outputs we call a $\mathfrak{v}$-task. The inputs and outputs in the $\mathfrak{v}$-task are ‘fit’ or ‘correct’, the rest are not. Given these inputs, the system is fit if it embodies a statement that constrains behaviour to correct outputs only. We call statements that constrain behaviour policies. Policies constrain outputs given inputs, as in the computer analogy. Just as many different neural topologies can have the same functionality, many different policies can imply the same set of outputs given a particular set of inputs. A system learns if it can construct a policy from historical data which constrains behaviour to correct outputs (generalises beyond historical data), and once again the optimal choice of policy is the weakest. For brevity, when a system optimises for weak policies, we say it is engaged in weak policy optimisation (WPO).
The above vocabulary can describe a distributed system. For example, an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$. The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [16] that, in its most basic form, differentiates observation from intervention in the manner of reafference [11, 30].
$\alpha_{1}$$\alpha_{2}$$\omega$
A policy is like an identity described in the MCA literature [45]. The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks (see proofs 1, 2 or [34]).
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$.
A collective of cells remains cells regardless of whether we deem it to be an organ. The parts are still there. Formally the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as the weakness proxy [34, proposition 1,2] facilitates causal learning at a single scale [16], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task. However, this still just treats everything as a collective of cells.
3.2 The Causal Chain of Abstraction
From the enactive perspective, a human is something cells do, rather than cells are.
To further clarify, we call a system complex when its behaviour is something we cannot easily comprehend or predict. However, just because we cannot comprehend and predict all of a system’s behaviour does not mean we cannot predict some of it. We suggest that so-called emergent properties are merely those aspects of a complex system’s behaviour which we can comprehend and predict. Aspects and patterns within patterns that remain comprehensible. For example, the overall pattern of a fractal like the Barnsley Fern [53] can be obvious, while the details elude comprehension.
We now illustrate how behaviour forms a causal chain of layers of increasingly abstract behaviour, from cells to organs to organisms to the behaviour of those organisms (specifically human abstractions). We’ll then show how it is possible to ’intervene’ in that causal chain to disconnect more abstract behaviour from the less abstract behaviour that gave rise to it. We show this reduces “utility” (def. 7) for tasks at higher levels, meaning reduced capacity for WPO. This is because the resulting distribution of control is altered when we intervene in the causal chain.
Just as an organ boils down to a collection of cells, the collective of cells as a whole is an abstraction layer in which “organ” runs, like software. However, software is just a state of hardware [12]. Put another way, “organ” is a classification we ascribe to the the behaviour of a cellular collective as a whole.
As stated in the introduction, we’re interested in how MCAs relate to abstraction layers in a computer. x86 is a classification we ascribe to the product of human behaviour, a CPU. It is a higher level behaviour of an MCA that includes humans. To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels. To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their collective behaviour, as a unified whole.
A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions). We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must more tightly constrain what is considered to be “correct” behaviour if that higher level is to exist in a meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Theorem 3 shows that the greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be. Intuitively, this just means adaptability later in the cause chain implies adaptability earlier in the chain.
Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
$cells$$humans$$abstractions$
When we create an abstraction layer like the ISA of a computer, we create a machine that embodies a narrow subset of our more abstract behaviours. On top of that ISA it can mimic all behaviours at higher levels of abstraction, because it is Turing complete. However, the ISA itself is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. No graceful degradation. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain and creates inefficiencies (which is why, for example, there is demand for application specific chipsets and field programmable gate arrays).
$humans$$ISA$$abstractions$
When we say the ISA is static, we mean there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. What proof 3 demonstrates is that adaptability at higher levels of abstraction demands adaptability at lower levels. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [31], because bounded systems can contain only finite information [54]. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
4 Drawbacks of Self Organisation
We have proposed and explored some advantages of multilayer-causal-learning. Are there disadvantages? Do such systems also inherit the flaws of biological systems?
Yes, they inherit a state analogous to cancer. In a biological system, when cells are isolated from the informational structure of their collective (i.e. an organ), they lose their high level identity. When that happens, they independently pursue their own lower level goals [55, 56]. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. This is what is commonly known as cancer.
In the context of our formalism, a state analogous to cancer occurs when a system is ’over-constrained’ by the abstraction layer in which it exists. To illustrate what we mean; the more difficult the task the more constrained the set of correct policies [42], meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. Collectives that exhibit ‘easier’ high level goal directed behaviour in that sense have more correct high level policies to choose from. They are less tightly constrained. If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder” in the sense of there existing fewer correct policies181818i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$. If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, then the only way the system can continue to function is for part of the collective to “splinter” or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind. In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to “selective forgetting” [16] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, we would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be under-constrained to ensure graceful degradation or “sloppy fitness” [57], so it can maintain core functionality in as wide a variety of unanticipated circumstances as possible. An under-constrained system may be flexible enough to cope with unexpected inputs that more tightly restrict the set of correct policies. For example, if environment conditions are hostile or “difficult”, that would constitute an input that constrains the space of correct policies. The more difficult the environment, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Conversely, if control of a system is distributed among its parts at the lowest levels (i.e. a swarm rather than a central controller), then the destruction of its parts should not inhibit higher level function. There would be little reason for the parts of a system to splinter, as that would not greatly increase the number of available correct policies.
5 Conclusions and Future Research
We have formalised MCAs as a causal chain, and shown that the capacity for WPO later in the chain depends upon WPO earlier in the chain. This seems to indicate MCA-like designs could result in more adaptable and reliable AI in future. Indeed, the layered structure of artificial neural networks resembles exactly the sort of structure we have described, albeit constrained by the static abstraction layers in which an artificial neural network runs and without WPO. This may explain something of their relative success as a machine learning method. Future research might consider how hardware could be designed in a more distributed, multiscale fashion in order to facilitate MCL by fine tuning abstraction layers. In the context of the current hardware, adaptive abstraction layers might be considered, to facilitate greater efficiency.
Beyond hardware, cyberphysical systems in general could be viewed through the lenses of abstraction and distribution we have described here, to ensure they are robust. We might consider how states analogous to cancer could arise in human organisational structures, or how power may be distributed throughout an organisation to ensure MCL. Free-market trade is a great example of how loose constraints have already facilitated adaptability in large collectives, and which is already being explored in the broader context of MCAs [58]. Likewise, the delegation of greater control has been shown to improve the effectiveness of military organisations [59]. Our claim boils down to a principle of design like an epistemological razor [34]; systems should be no more constrained than absolutely necessary.
One concrete example of how this might be applied in future is the regulation of AI. The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [60]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [61]. This involves constructing advanced general AI of level ASL-4 [62] [^19] Meaning an artificial general intelligence [20] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI. in a cyberphysical system. That system seems intended to safeguard against threats from both within and without. The idea of a “rogue AI” from within this system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective. However if we view this system as MCA, then tight regulations may over-constrain the system trigger the cancerous state they are meant to prevent.
This also might be applied to problems of a philosophical nature. Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole). Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [16]. By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. This relates to the hard problem of consciousness [63]. All abstractions constructed by living things are “goal directed” in the sense that they are consequence of an evolutionary process [30]. In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them. It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements). Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
In summary, we have formally described biological self organisation across distribution of scale and levels of abstraction, and shown adaptability at higher levels of abstraction depends upon adaptability at lower levels. We have described how a failure state analogous to cancer can occur when a system is too tightly constrained by the abstraction layer in which it exists (i.e. the task it faces is impossible). Though our result is of a purely theoretical nature, it illustrates how we may design more robust systems, and lays a foundation for future empirical research.